# fMoE: Fine-Grained Expert Offloading for Large Mixture-of-Experts Serving

### v1

@@ Abstract
Large Language Models (LLMs) have gained immense success in revolutionizing various applications, including content generation, search and recommendation, and AI-assisted operation. To reduce high training costs, Mixture-of-Experts (MoE) architecture has become a popular backbone for modern LLMs. However, despite the benefits, serving MoE-based LLMs experience severe memory inefficiency due to sparsely activated experts. Recent studies propose to offload inactive experts from GPU memory to CPU memory to improve the serving efficiency of MoE models. However, they either incur high inference latency or high model memory footprints due to coarse-grained designs. To tame the latency-memory trade-off in MoE serving, we present fMoE, a fine-grained expert offloading system for MoE serving that achieves low inference latency with memory efficiency. We design fMoE to extract fine-grained expert selection patterns from MoE models and semantic hints from input prompts to efficiently guide expert prefetching, caching, and offloading decisions. fMoE is prototyped on top of HuggingFace Transformers and deployed on a six-GPU testbed. Experiments with open-source MoE models and real-world workloads show that fMoE reduces inference latency by 47% and improves expert hit rate by 36% over state-of-the-art solutions.
1 Introduction
@@
Large Language Models have achieved remarkable success in advancing Natural Language Processing (NLP) research and transforming various applications, including content generation [11, 2, 6, 40], search and recommendation [30, 59], and AI-assisted operations [34, 29, 22]. Given the high training costs, modern LLMs have returned to Mixture-of-Experts (MoE) architectures [21, 46, 57, 53, 10, 1] as their backbone implementations. Inside MoE models, each MoE layer comprises a gating network and a collection of experts, with only a subset of experts being activated during computation. This sparse activation mechanism significantly reduces the number of floating point operations, enabling MoE-based LLMs to achieve substantially lower training costs compared to dense LLMs [10, 57, 21].
@@
Despite the computational efficiency, MoE models exhibit substantial memory inefficiency during the serving phase. Though certain model parameters remain inactive during inference, they must still reside in GPU memory to allow for potential future activation. Expert offloading [54, 47, 16, 4] has emerged as a promising strategy to address this issue, which predicts inactive experts and transfers them to CPU memory while retaining only the necessary experts in GPU memory, reducing the overall model memory footprint.
@@
However, existing expert offloading solutions struggle to effectively balance the latency-memory trade-off in MoE serving. These approaches either suffer from high inference latency [4, 47] or incur substantial model memory footprints [16, 54]. The key reason is that existing works track expert patterns and manage experts in coarse granularity. They fail to accurately identify and retain only the necessary experts in GPU memory during inference, resulting in frequent and costly on-demand expert loading [47], which severely degrades serving performance.
@@
In this paper, we propose fMoE, a fine-grained expert offloading system that tames the latency-memory trade-off in MoE serving. To track and analyze MoE models’ expert selection behaviors in fine granularity, we propose a new data structure called expert map, which records the iteration-level probability distributions output by the gate network. fMoE uses historical expert maps for comparing expert trajectory similarity to guide offloading.111In this paper, “trajectory” is defined as the collection of probability distributions over experts observed through layers. Apart from the expert map, fMoE is designed to track fine-grained input semantic embeddings from individual request prompts processed by the MoE model. Given the collected semantic-based and trajectory-based information, fMoE carefully searches the most accurate expert map for guiding expert prefetching, caching, and offloading through inference iterations. In summary, we make the following contributions:
@@
We design fMoE, a fine-grained expert offloading system that achieves low inference latency while reducing model memory footprints.
@@
We propose a new data structure, expert map, that tracks fine-grained expert selection behaviors of MoE models. fMoE leverages input semantic embeddings to augment the expert map search for guiding expert offloading.
@@
We prototype fMoE on top of HuggingFace Transformers [52] and deploy it on a six-GPU testbed. Extensive experiments with open-source MoE models and real-world workloads show that fMoE reduces inference latency by 47% and improves expert hit rate by 36% compared to state-of-the-art solutions.
2 Background and Motivation
2.1 LLM Serving
@@
Unlike traditional Deep Learning (DL) model inference, Large Language Model (LLM) serving consists of two consecutive stages: prefill and decode. Figure 3 illustrates the two stages when an LLM performs inference for a request prompt. In the prefill stage, the LLM first computes the intermediate key-value (KV) states of the prompt tokens, prefills the KV cache [24, 32, 27, 61, 3], and then generates the first answer token. In the decode stage, the LLM sequentially generates the answer to the prompt token-by-token in an auto-regressive manner, where tokens generated previously are used for generating the next token.
@@
The two stages have their own unique characteristics. The prefill stage only requires one iteration, processing all tokens in parallel and generating the first answer token. The decode stage spans several iterations, generating one token per iteration until the answer is completed. Due to the different characteristics of the two stages, recent studies [38, 61] have identified that the prefill stage is compute-bounded, while the decode stage is considered memory-bounded. Therefore, people typically quantify the serving performance of LLM two stages using different metrics. For the prefill stage, Time-To-First-Token (TTFT) is commonly employed, which measures the latency from receiving the user request until generating the first answer token. For the decode stage, Tokens-Per-Second (TPS) or Time-Per-Output-Token (TPOT) is used to measure the generation rate of LLM serving.
2.2 MoE-based LLM Serving
@@
By integrating MoE layers in Transformer blocks [51], MoE architectures [58] have emerged as a popular backbone for modern LLMs, such as Mixtral [21], Snowflake Arctic [46], and DeepSeek-MoE [10]. Figure 3 illustrates MoE-based LLMs’ typical structures, where feed-forward network (FFN) modules are replaced by MoE layers. Each MoE layer consists of a gate network and a set of expert networks. Inside each Transformer block, the self-attention module first calculates the attentions [51] based on input hidden states, and then the gate network determines which expert(s) to activate for computing the output representations. Compared to traditional dense LLMs, MoE-based LLMs only activate a subset of parameters during training and inference, reducing computational overhead while delivering superior generation performance compared to dense LLMs with a comparable number of parameters [21, 46, 10, 53, 57, 1].
@@
Despite the benefits of saving training computations, MoE-based LLM serving still suffers from GPU memory inefficiency as MoE inference requires loading all model parameters into GPU memory, including those inactive experts. For example, Mixtral-87B [21] and DeepSeek-MoE [10] have 72% and 83% inactive parameters during inference due to the sparsity of expert activation in MoE, leading to low memory efficiency and serving throughput. Therefore, to efficiently serve large MoE models, we must seek a solution to the memory inefficiency inherited from MoE architecture.
2.3 Latency-Memory Trade-Off
@@
Recently, a few studies have been proposed to improve MoE-based LLM serving efficiency. Figure 2 describes the design space in MoE serving. Existing major studies can be categorized into two types: Lossy serving applies compression [39], pruning [26], and quantization [23] techniques to the original MoE models to reduce the serving memory requirements. However, this line of work achieves serving efficiency by sacrificing the generation quality. Lossless serving focuses on offloading model weights (parameters [4, 36] or experts [16, 47, 54]) that are sparsely utilized in temporal or spatial patterns from GPU memory to CPU memory, aiming to preserve reasonable inference latency. Specifically, expert offloading seeks to predict the activation of experts in advance, prefetching or caching only the necessary experts in GPU memory during inference. We opt for lossless serving to design fMoE because this line of methods avoids modifying models, hence assuring generation quality.
@@
However, existing offloading solutions cannot achieve an optimal spot in the latency-memory trade-off when serving MoE-based LLMs. Figure 1(b) compares the performance (i.e., inference latency and memory footprint) of existing state-of-the-art (SOTA) offloading solutions, which either provide low inference latency but suffer from large memory footprint (e.g., No-offload and MoE-Infinity [54]), or vice versa (e.g., ProMoE [47], Mixtral-Offloading [16], and DeepSpeed-Inference [4]).
@@
The key reason behind this dilemma is that MoE-based decoder-only LLMs have balanced expert routing [47], leaving existing solutions hard to find effective patterns for guiding expert offloading. Existing research has identified two main reasons for this dilemma: First, most MoE-based LLMs are decoder-only architectures, which exhibit uniform expert activation patterns and low expert access skewness compared to encoder-decoder MoE LLMs [47, 18]. Second, recent MoE-based LLMs are trained with a unique load balancing loss [21, 46, 53, 10, 1], which enforces gate networks to balance the tokens routed to each expert within the same MoE layer, ensuring no experts are trivial throughout training. This balanced routing diminishes the predictability of expert patterns, thus making existing solutions ineffective.
2.4 Existing MoE Offloading Solutions
@@
Existing expert offloading approaches [16, 54] rely on coarse-grained expert patterns, which are inefficient for guiding offloading. We define coarse-grained information as the expert patterns collected at the request level, where information is aggregated over multiple iterations of a request prompt. For example, MoE-Infinity [54] tracks request-level expert activations. Fine-grained information is defined as the expert patterns observed separately during each inference iteration. Figure 3(a) shows examples of coarse-grained and fine-grained expert activation heatmaps for Mixtral-87B [21]. The heatmap records the expert activations across 32 MoE layers, where each layer contains eight experts and activates two experts out of eight to compute representations. While fine-grained (iteration-level) heatmaps show clear expert activation patterns, the aggregated coarse-grained (request-level) heatmap diminishes predictability.
@@
To demonstrate this point, we analyze the Shannon entropy [44] of expert activations per MoE layer for three popular MoE models. Entropy is an essential metric to quantify the uncertainty and unpredictability of variables in information theory. A balanced expert activation pattern (e.g., probability distribution of four experts) results in a high entropy, which indicates the pattern is more unpredictable and hard to select experts. Figure 3(b) presents the mean entropy computed per layer for three MoE models (Mixtral-87B [21], Qwen1.5-MoE [57], and Phi-3.5-MoE [1]) across two real-world datasets (LMSYS-Chat-1M [60] and ShareGPT [45]). Coarse-grained expert patterns have significantly higher entropy than fine-grained patterns, meaning that expert patterns in coarse granularity can be less effective for predictions. Figure 3(c) shows the mean entropy per layer through inference iterations. While the entropy is low at the beginning of inference, it gradually increases through iterations due to aggregating expert activation information, thus becoming more unpredictable.
@@
2.5 Problems of Coarse-Grained Offloading
@@
1) Insufficient latency-memory trade-off. Existing solutions prefetch and offload experts in coarse granularity, either heavily focusing on reducing inference latency but incurring large memory footprint [54] or reducing memory footprint but severely increasing inference latency [4, 16].
2) Low expert hit rates. Existing solutions employ coarse-grained expert pattern tracking methods (e.g., Expert Activation Matrix in MoE-Infinity [54]), which produce ineffective expert patterns for guiding offloading decisions, leading to low expert hit rates and high inference latency.
3) Ignorance of MoE models’ and prompts’ heterogeneity. Existing solutions largely ignore the unique characteristics of different MoE models and input prompts and serve them in a one-fits-all manner [4, 16, 47, 54], which omits opportunities for fine-grained optimizations adaptive to heterogeneous models and prompts in MoE serving.
Figure 4 shows the expert hit rates of serving three popular MoE-based LLMs, Mixtral-87B [21], Qwen1.5-MoE [57], and Phi-3.5-MoE [1] using LMSYS-Chat-1M dataset [60] with coarse-grained and fine-grained expert offloading designs at different prefetch distances, respectively. Prefetch distance refers to the number of layers ahead that a prefetch instruction is issued before the target layer activates its experts. By leveraging fine-grained expert offloading, we can achieve significantly higher expert hit rates over coarse-grained methods and preserve better performance by adapting to varying prefetch distances.
3 fMoE’s Overview
3.1 Objectives and Challenges
fMoE is designed to achieve the following three goals:
Memory-efficient MoE serving with minimal inference latency. We have demonstrated that existing expert offloading solutions [16, 47, 54] fail to tame the latency-memory trade-off in MoE serving (§2.3). We aim to achieve both low memory footprint and inference latency by proposing fine-grained expert offloading.
Minimize expert miss due to mispredictions in expert prefetching. Expert prefetching, involving future expert activation predictions, is an essential step in expert offloading solutions. However, a recent study [47] has shown that expert miss due to mispredictions can cause high on-demand expert loading delay in inference. We should minimize expert miss and mitigate mispredictions in expert offloading.
Adapt to heterogeneous MoE models and prompts. MoE inference can serve heterogeneous models [21, 10, 57, 46, 53] with varying prompts [60, 45] in real-world scenarios. While existing solutions handle different models and prompts with a one-fits-all design, we should design our expert offloading to adapt to the heterogeneity in MoE serving.

### v2

# Taming Latency-Memory Trade-Off in MoE-Based LLM Serving via Fine-Grained Expert Offloading

@@ Abstract.
Large Language Models (LLMs) have gained immense success in revolutionizing various applications, including content generation, search and recommendation, and AI-assisted operations. To reduce high training costs, Mixture-of-Experts (MoE) architecture has become a popular backbone for modern LLMs. However, despite the benefits, serving MoE-based LLMs experience severe memory inefficiency due to sparsely activated experts. Recent studies propose to offload inactive experts from GPU memory to CPU memory to improve the serving efficiency of MoE models. However, they either incur high inference latency or high model memory footprints due to coarse-grained designs.
To tame the latency-memory trade-off in MoE serving, we present FineMoE, a fine-grained expert offloading system for MoE serving that achieves low inference latency with memory efficiency. We design FineMoE to extract fine-grained expert selection patterns from MoE models and semantic hints from input prompts to efficiently guide expert prefetching, caching, and offloading decisions. FineMoE is prototyped on top of HuggingFace Transformers and deployed on a six-GPU testbed. Experiments with open-source MoE models and real-world workloads show that FineMoE reduces inference latency by 47% and improves expert hit rate by 39% over state-of-the-art solutions.
1. Introduction
@@
Large Language Models have achieved remarkable success in advancing Natural Language Processing (NLP) research and transforming various applications, including content generation (Dai et al., 2024b; Achiam et al., 2023; Brown et al., 2020; Radford et al., 2019), search and recommendation (Lin et al., 2024; Zhao et al., 2024), and AI-assisted operations (Nam et al., 2024; Li et al., 2024; Jiang et al., 2024a). Given the high training costs, modern LLMs have returned to Mixture-of-Experts (MoE) architectures (Jiang et al., 2024b; Snowflake, 2024; Yang et al., 2024; xAI, 2023; Dai et al., 2024a; Abdin et al., 2024) as their backbone implementations. Inside MoE models, each MoE layer comprises a gating network and a collection of experts, with only a subset of experts being activated during computation. This sparse activation mechanism significantly reduces the number of floating point operations, enabling MoE-based LLMs to achieve substantially lower training costs compared to dense LLMs (Dai et al., 2024a; Yang et al., 2024; Jiang et al., 2024b).
@@
Despite the computational efficiency, MoE models exhibit substantial memory inefficiency during the serving phase. Though certain model parameters remain inactive during inference, they must still reside in GPU memory to allow for potential future activation. Expert offloading (Xue et al., 2024; Song et al., 2024; Eliseev and Mazur, 2023; Aminabadi et al., 2022) has emerged as a promising strategy to address this issue, which predicts inactive experts and transfers them to CPU memory while retaining only the necessary experts in GPU memory, reducing the overall model memory footprint.
@@
However, existing expert offloading solutions struggle to effectively balance the latency-memory trade-off in MoE serving. These approaches either suffer from high inference latency (Aminabadi et al., 2022; Song et al., 2024) or incur substantial model memory footprints (Eliseev and Mazur, 2023; Xue et al., 2024). The key reason is that existing works track expert patterns and manage experts in coarse granularity. They fail to accurately identify and retain only the necessary experts in GPU memory during inference, resulting in frequent and costly on-demand expert loading (Song et al., 2024) and performance degradation.
@@
In this paper, we propose FineMoE, a fine-grained expert offloading system that tames the latency-memory trade-off in MoE serving. To track and analyze MoE models’ expert selection behaviors in fine granularity, we propose a new data structure called expert map, which records the iteration-level probability distributions output by the gate network. FineMoE uses historical expert maps for comparing expert trajectory similarity to guide offloading.111In this paper, “trajectory” is defined as the collection of probability distributions over experts observed through layers. Apart from the expert map, FineMoE is designed to track fine-grained input semantic embeddings from individual request prompts processed by the MoE model. Given the collected semantic-based and trajectory-based information, FineMoE carefully searches the most accurate expert map for guiding expert prefetching, caching, and offloading through inference iterations. In summary, we make the following contributions:
@@
We design FineMoE, a fine-grained expert offloading system that achieves low inference latency while reducing model memory footprints.
@@
We propose a new data structure, expert map, that tracks fine-grained expert selection behaviors of MoE models. FineMoE leverages input semantic embeddings to augment the expert map search to guide expert offloading.
@@
We prototype FineMoE on top of HuggingFace Transformers (Wolf et al., 2020) and deploy it on a six-GPU testbed. Extensive experiments with open-source MoE models and real-world workloads show that FineMoE reduces inference latency by 47% and improves expert hit rate by 39% compared to state-of-the-art solutions.
2. Background and Motivation
2.1. LLM Serving
@@
Unlike traditional Deep Learning (DL) model inference, Large Language Model (LLM) serving consists of two consecutive stages: prefill and decode. Figure 1(a) illustrates the two stages when an LLM performs inference for a request prompt. In the prefill stage, the LLM first computes the intermediate key-value (KV) states of the prompt tokens, prefills the KV cache (Kwon et al., 2023; Liu et al., 2024b; Lee et al., 2024a; Zhong et al., 2024; Agrawal et al., 2024), and then generates the first answer token. In the decode stage, the LLM sequentially generates the answer to the prompt token-by-token in an auto-regressive manner, where tokens generated previously are used for generating the next token.
@@
The two stages have their own unique characteristics. The prefill stage only requires one iteration222An iteration refers to a single step in auto-regressive inference that generates one new token. The iteration time denotes the end-to-end latency of this step., processing all tokens in parallel and generating the first answer token. The decode stage spans several iterations, generating one token per iteration until the answer is completed. Due to the different characteristics of the two stages, recent studies (Patel et al., 2024; Zhong et al., 2024) have identified that the prefill stage is compute-bounded, while the decode stage is considered memory-bounded. Therefore, people typically quantify the serving performance of LLM two stages using different metrics. For the prefill stage, Time-To-First-Token (TTFT) is commonly employed, which measures the latency from receiving the user request until generating the first answer token. For the decode stage, Tokens-Per-Second (TPS) or Time-Per-Output-Token (TPOT) is used to measure the generation rate of LLM serving.
2.2. MoE-based LLM Serving
@@
By integrating MoE layers in Transformer blocks (Vaswani, 2017), MoE architectures (Yuksel et al., 2012) have emerged as a popular backbone for modern LLMs, such as Mixtral (Jiang et al., 2024b), Snowflake Arctic (Snowflake, 2024), and DeepSeek-MoE (Dai et al., 2024a). Figure 1(a) illustrates MoE-based LLMs’ typical structures, where feed-forward network (FFN) modules are replaced by MoE layers.333For simplicity, we only show the process of one single request prompt. Each MoE layer consists of a gate network and a set of expert networks. Inside each Transformer block, the self-attention module first calculates the attentions (Vaswani, 2017) based on input hidden states, and then the gate network determines which expert(s) to activate for computing the output representations. Compared to traditional dense LLMs, MoE-based LLMs only activate a subset of parameters during training and inference, reducing computational overhead while delivering superior generation performance compared to dense LLMs with a comparable number of parameters (Jiang et al., 2024b; Snowflake, 2024; Dai et al., 2024a; xAI, 2023; Yang et al., 2024; Abdin et al., 2024).
@@
Despite the benefits of saving training computations, MoE-based LLM serving still suffers from GPU memory inefficiency as MoE inference requires loading all model parameters into GPU memory, including those inactive experts. Table 1 characterizes three popular MoE models: Mixtral-87B (Jiang et al., 2024b), Qwen1.5-MoE (Yang et al., 2024), and Phi-3.5-MoE (Abdin et al., 2024). During inference, they exhibit 72%, 81%, and 84% inactive parameters, respectively, due to the sparsity of expert activation in MoE. This corresponds to 67, 23, and 70 GB of inactive GPU memory, resulting in low memory efficiency and serving throughput. Therefore, to efficiently serve large MoE models, we must seek a solution to the memory inefficiency inherited from MoE architecture.
2.3. Latency-Memory Trade-Off
@@
Recently, a few studies have been proposed to improve MoE-based LLM serving efficiency. Figure 2 describes the design space in MoE serving. Existing major studies can be categorized into two types: Lossy serving applies compression (Pingzhi Li, Zhenyu Zhang, Prateek Yadav, Yi-Lin Sung, Yu Cheng, Mohit Bansal, Tianlong Chen, 2024), pruning (Lee et al., 2024b), and quantization (Kim et al., 2023) techniques to the original MoE models to reduce the serving memory requirements. However, this line of work achieves serving efficiency by sacrificing the generation quality. Lossless serving focuses on offloading model weights (parameters (Aminabadi et al., 2022; Ollama, 2024) or experts (Eliseev and Mazur, 2023; Song et al., 2024; Xue et al., 2024)) that are sparsely utilized in temporal or spatial patterns from GPU memory to CPU memory, aiming to preserve reasonable inference latency. Specifically, expert offloading seeks to predict the activation of experts in advance, prefetching or caching only the necessary experts in GPU memory during inference. We opt for lossless serving to design FineMoE because this line of methods avoids modifying models, hence assuring generation quality.
@@
However, existing offloading solutions cannot achieve an optimal spot in the latency-memory trade-off when serving MoE-based LLMs. Figure 1(b) compares the performance (i.e., inference latency and memory footprint) of existing state-of-the-art (SOTA) offloading solutions, which either provide low inference latency but suffer from large memory footprint (e.g., No-offload and MoE-Infinity (Xue et al., 2024)), or vice versa (e.g., ProMoE (Song et al., 2024), Mixtral-Offloading (Eliseev and Mazur, 2023), and DeepSpeed-Inference (Aminabadi et al., 2022)).
@@
The key reason behind this dilemma is that MoE-based decoder-only LLMs have balanced expert routing (Song et al., 2024), leaving existing solutions hard to find effective patterns for guiding expert offloading. Existing research has identified two main reasons for this dilemma: First, most MoE-based LLMs are decoder-only architectures, which exhibit uniform expert activation patterns and low expert access skewness compared to encoder-decoder MoE LLMs (Song et al., 2024; Gupta et al., 2024). Second, recent MoE-based LLMs employ a load-balancing loss (Jiang et al., 2024b; Snowflake, 2024; xAI, 2023; Dai et al., 2024a; Abdin et al., 2024), which encourages the gate network to distribute tokens more uniformly across experts within each MoE layer, making expert usage more balanced during training. This balanced routing diminishes the predictability of expert patterns, thus making existing solutions ineffective.
2.4. Existing MoE Offloading Solutions
@@
Existing expert offloading approaches (Eliseev and Mazur, 2023; Xue et al., 2024) rely on coarse-grained expert patterns, which are inefficient for guiding offloading. We define coarse-grained information as the expert patterns collected at the request level, where information is aggregated over multiple iterations of a request prompt. For example, MoE-Infinity (Xue et al., 2024) tracks request-level expert activations. Fine-grained information is defined as the expert patterns observed separately during each inference iteration. Figure 3(a) shows examples of coarse-grained and fine-grained expert activation heatmaps for Mixtral-87B (Jiang et al., 2024b). The heatmap records the expert activations across 32 MoE layers, where each layer contains eight experts and activates two experts out of eight to compute representations. While fine-grained (iteration-level) heatmaps show clear expert activation patterns, the aggregated coarse-grained (request-level) heatmap diminishes predictability.
@@
To demonstrate this point, we analyze the Shannon entropy (Shannon, 1948) of expert activations per MoE layer for three popular MoE models. Entropy is an essential metric to quantify the uncertainty and unpredictability of variables in information theory. A balanced expert activation pattern (e.g., probability distribution of four experts) results in a high entropy, which indicates the pattern is less predictable and harder to select experts. Figure 3(b) presents the mean entropy computed per layer for three MoE models (Mixtral-87B (Jiang et al., 2024b), Qwen1.5-MoE (Yang et al., 2024), and Phi-3.5-MoE (Abdin et al., 2024)) across two realistic datasets LMSYS-Chat-1M (Zheng et al., 2023) and ShareGPT (ShareGPT, 2022). Coarse-grained expert patterns have significantly higher entropy than fine-grained patterns, meaning that expert patterns in coarse granularity can be less effective for predictions. Figure 3(c) shows the mean entropy per layer when aggregating expert patterns across inference iterations, where expert selection becomes increasingly unpredictable as generation progresses. Qwen1.5-MoE reaches a higher entropy plateau due to its larger expert selection space (60 experts × 24 layers). Similarly, Phi-3.5-MoE (16 × 32) exhibits higher entropy than Mixtral-87B (8 × 32). After about ten iterations, expert patterns become blurred and the entropy plateaus, indicating that further iterations contribute only marginal additional unpredictability. While entropy is low at the beginning of inference, it gradually increases with iterations as more expert activation information is aggregated, thereby becoming more unpredictable.
@@
2.5. Problems of Coarse-Grained Offloading
@@
1) Insufficient latency-memory trade-off. Existing solutions prefetch and offload experts in coarse granularity, either heavily focusing on reducing inference latency but incurring large memory footprint (Xue et al., 2024) or reducing memory footprint but severely increasing inference latency (Aminabadi et al., 2022; Eliseev and Mazur, 2023).
2) Low expert hit rates. Existing solutions employ coarse-grained expert pattern tracking methods (e.g., Expert Activation Matrix in MoE-Infinity (Xue et al., 2024)), which produce ineffective expert patterns for guiding offloading decisions, leading to low expert hit rates and high inference latency.
3) Ignorance of MoE models’ and prompts’ heterogeneity. Existing solutions largely ignore the unique characteristics of different MoE models and input prompts and serve them in a one-fits-all manner (Aminabadi et al., 2022; Eliseev and Mazur, 2023; Song et al., 2024; Xue et al., 2024), which omits opportunities for fine-grained optimizations adaptive to heterogeneous models and prompts in MoE serving.
Figure 4 shows the expert hit rates of serving three popular MoE-based LLMs, Mixtral-87B (Jiang et al., 2024b), Qwen1.5-MoE (Yang et al., 2024), and Phi-3.5-MoE (Abdin et al., 2024) using LMSYS-Chat-1M dataset (Zheng et al., 2023) with coarse-grained and fine-grained expert offloading designs at different prefetch distances, respectively. Prefetch distance refers to the number of layers ahead that a prefetch instruction is issued before the target layer activates its experts. By leveraging fine-grained expert offloading, we can achieve significantly higher expert hit rates over coarse-grained methods and preserve better performance by adapting to varying prefetch distances.
3. FineMoE’s Overview
3.1. Objectives and Challenges
FineMoE is designed to achieve the following three goals:
Memory-efficient MoE serving with minimal inference latency. We have demonstrated that existing expert offloading solutions (Eliseev and Mazur, 2023; Song et al., 2024; Xue et al., 2024) fail to tame the latency-memory trade-off in MoE serving (§2.3). We aim to achieve both low memory footprint and inference latency by proposing fine-grained expert offloading.
Minimize expert miss due to mispredictions in expert prefetching. Expert prefetching, involving future expert activation predictions, is an essential step in expert offloading solutions. However, a recent study (Song et al., 2024) has shown that expert miss due to mispredictions can cause high on-demand expert loading delay in inference. We should minimize expert miss and mitigate mispredictions in expert offloading.
Adapt to heterogeneous MoE models and prompts. MoE inference can serve heterogeneous models (Jiang et al., 2024b; Dai et al., 2024a; Yang et al., 2024; Snowflake, 2024; xAI, 2023) with varying prompts (Zheng et al., 2023; ShareGPT, 2022) in real-world scenarios. While existing solutions handle different models and prompts with a one-fits-all design, we should design our expert offloading to adapt to the heterogeneity in MoE serving.